OpenStudy (anonymous):

Does anyone know how to set the problem up? In the savannah, a sinkhole opened up catching some ostrich and cheetah. There was a total of 13 animals to be rescued. If 44 individual feet were pulled from the muck, how many ostrich and how many cheetahs were rescued?

OpenStudy (phi):

you have to count feet, knowing that ostriches have 2 feet and cheetah's have 4 feet first step: set up an equation for the number of feet if you have x ostriches and y cheetahs

OpenStudy (anonymous):

9 cheetah and 4 ostrich

OpenStudy (anonymous):

x(2)+y(4)=44?

OpenStudy (ganpat):

and one more equation.. can u guess ??

OpenStudy (phi):

yes now an equation for the number of animals if you have x ostriches and y cheetahs.

OpenStudy (anonymous):

hmmmm. i honestly have no idea

OpenStudy (anonymous):

x+y gives 13

OpenStudy (phi):

think about it: you have x ostriches and y cheetahs.

OpenStudy (anonymous):

but i just made an equation with x and y?

OpenStudy (phi):

that's ok, you can make another one (if fact, that is the point) your first one was about feet. this next one is about heads (everybody has only one head)

OpenStudy (anonymous):

oh so x+y=44?

OpenStudy (phi):

44 is the number of feet. look at the problem.

OpenStudy (anonymous):

x+y=13

OpenStudy (phi):

does that make sense?

OpenStudy (anonymous):

yes! thankyou. no could tell me how to set it up? would i put the first equation over the second one or solve them separately?

OpenStudy (anonymous):

*now

OpenStudy (phi):

we have 2 equations and 2 unknowns (we always need as many equations as unknowns) 2x+4y=44 x+y= 13 I would multiply the bottom equation by -2 (that means multiply every term on both sides of = by -2) what do you get?

OpenStudy (phi):

then add the two equations.

OpenStudy (anonymous):

2x=4y=44 -x^2+-y^2= -26?

OpenStudy (phi):

multiply like this: put a -2* in front of every term what do you get?

OpenStudy (anonymous):

-2x+-2y+ -26?

OpenStudy (phi):

ok, but you left out the = sign so what we now have is 2x+4y=44 -2x+-2y= -26

OpenStudy (phi):

now add the 2 equations. the easy way is just write down all the terms on one line on the left side of the = and on the right side

OpenStudy (anonymous):

I added and got 18

OpenStudy (phi):

that is the right side. do the same thing for the left side. and you should write down the whole equation (less chance of making a mistake)

OpenStudy (anonymous):

the right side?

OpenStudy (phi):

like this: 2x+ 4y= 44 -2x+-2y= -26 ------------- 2x-2x+4y+ -2y= 44-26 now simplify. the right side (as you know) is 18 can you simplify the left side?

OpenStudy (anonymous):

x cancels out i think and then you have 2y=18 /2 /2 y=9 but whatdoes x equal?

OpenStudy (phi):

looks good. now go back to the original equations 2x+4y=44 x+y= 13 pick one of them (does not matter which, except the 2nd one looks simpler) replace y with 9

OpenStudy (anonymous):

x+y= 13 x+9=13 -9 -9 x=4 THANKYOUSO MUCH FOR ALL YOUR HELP :)

OpenStudy (phi):

yw. look over how we did it. do you see why we multiplied the 2nd equation by -2 before adding the 2 equations? (to make the x's cancel out) so one of the tricks is seeing that -2 works.